Problem 1)

The management of VITULLO Company is trying to decide whether to continue manufacturing a part or to buy it from an outside supplier. The part, called FIZBE, is a component of the company’s finished product.

         The following information was collected from the accounting records and production data for the year ending December 31, 2017.

  1. 5,000 units of FIZBE were produced in the machining department.
  2. Variable manufacturing costs applicable to the production each FIZBE unit were: direct materials $4.75, direct labor $4.60, Indirect labor $0.45, utilities $0.35.
  3. Fixed manufacturing costs applicable to the production of FIZBE were:

COST

DIRECT

ALLOCATED

Depreciation

$1,100

$900

Property Taxes

500

200

Insurance

900_____

600_____

$2500

$1700

All variable manufacturing and direct fixed costs will be eliminated if FIZBE is purchased. Allocated costs will have to be absorbed by other production departments.

  1. The lowest quotation for 5,000 FIZBE units from a supplier is $56,000.
  2. If FIZBE units are purchased, freight and inspection costs would be $0.30 per unit, and receiving costs totaling $500 per year would be incurred by the Machining Department.

Instructions

  1. Prepare an incremental analysis for FIZBE. Your analysis should have columns for (1) Make FIZBE, (2) Buy FIZBE, and (3) Net Income Increase/Decrease.
  2. Based on your analysis, what decision should management make?
  3. Would the decision be different if VITULLO Company has the opportunity to produce $6,000 of net income with the facilities currently being used to manufacture FIZBE? Show Computations.
  4. What nonfinancial factors should management consider in making its decision?

Problem 2)

Panda Corporation has four operating divisions. During the first quarter of 2014, the company reported aggregate income from operations of $129,000 and the divisional results shown below.

                                                                                  Division

       I____

    ___II____

      _III____

       IV____

Sales

$510,000

$400,000

$310,000

$170,000

Cost of Goods Sold

300,000

250,000

270,000

156,000

Selling and Administration Expenses

60,000

80,000

75,000

70,000

Income (loss) from operations

$150,000

$70,000

$(35,000)

$(56,000)

Analysis Reveals the following percentages of variable costs in each division.

                                                                                  Division

       I____

    ___II____

      _III____

       IV____

Cost of Goods Sold

70%

80%

70%

90%

Selling and Administration Expenses

40%

50%

60%

70%

Discontinuance of any division would save 50% of the fixed costs and expenses for that division.

Top management is very concerned about the unprofitable divisions (III and IV). Consensus is that one or both of the divisions should be discontinued.

Instructions

  1. Compute contribution margin for Divisions III and IV.
  2. Prepare an incremental analysis concerning the possible discontinuance of (1) Division III and (2) Division IV. What course of action do you recommend for each division?

Extra Credit

  1. Prepare a columnar condensed income statement for Panda Corporation, assuming Division IV is eliminated. (Use the CVP format). Division IV’s unavoidable fixed costs are allocated equally to the continuing divisions.
  2. Reconcile the total income from operations ($129,000) with the total income from operations without Division IV.
Answer #1

Problem 1 .

Make or Buy Fibze

Make Cost Incurred

Direct Material   = 5000*4.75

Direct Labour = 5000*4.60

Indirect Labour = 5000.0.45

Utilites = 5000*.35

Total Direct Cost = 5000* 10.15 = $ 50750 + 2500+1700 = $ 54950

Buy 5000 @5

  1. Make or Buy Fibze

Make Cost Incurred

Direct Material   = 5000*4.75

Direct Labour = 5000*4.60

Indirect Labour = 5000.0.45

Utilites = 5000*.35

Total Direct Cost = 5000* 10.15 = $ 50750 + 2500+1700 = $ 54950

Buy 5000 @56000 + freight $ 1500 +Receiving cost $ 500

Total Cost on buy would be $ 58000

  1. If bought it would result in incremental cost of $ 58000-53250 = 4750
  2. The management should continue to produce rather than procure 5000 units as buy would result in incremental cost of $ 4750
  3. The direct cost to manufacture is $ 10.65 and procurement cost is 11.2 , which is higher , hence it would be advisable to produce 6000 units rather than buy.
  4. The non financial factors that affect the decision of management are the capacity of production , the quality of product and the relations with the suppliers .

6000 + freight $ 1500 +Receiving cost $ 500

Total Cost on buy would be $ 58000

  1. If bought it would result in incremental cost of $ 58000-53250 = 4750
  2. The management should continue to produce rather than procure 5000 units as buy would result in incremental cost of $ 4750
  3. The direct cost to manufacture is $ 10.65 and procurement cost is 11.2 , which is higher , hence it would be advisable to produce 6000 units rather than buy.
  4. The non financial factors that affect the decision of management are the capacity of production , the quality of product and the relations with the suppliers .

Problem 2

i) Contribution Margin

Particulars I II III IV
Sales 510000 400000 310000 170000
VC 70 80 70 90
Selling Distribution 40 50 60 70
Cost of Goods Sold 300000 250000 270000 156000
SD 60000 80000 75000 70000
VC 234000 240000 234000 189400
Contribution 276000 160000 76000 -19400
Fixed Cost 126000 90000 111000 36600
Income 150000 70000 -35000 -56000

ii)

Division IV Division III
VC 189400 234000
Sales 170000 310000
Decrease in revenue -19400 76000

To Close down division IV would be good decision as it is generating negative contribution .

iii) if Division IV is eliminated , the income statement would be

Fixed Cost of Division IV would be 50% = 36600 *50/100 = 18300 to be allocated equally among three divisions

Sales 510000 400000 310000
VC 70 80 70
Selling Distribution 40 50 60
Cost of Goods Sold 300000 250000 270000
SD 60000 80000 75000
VC 234000 240000 234000
Contibution 276000 160000 76000
Fixed Cost 126000 90000 111000
Allocated FD 6100 6100 6100
Net Income 143900 63900 -41100

iv) The total income from operations without Division IV would be 166700 ( 143900+63900-41100)